# Bringing Back the Context: Camera Trap Species Identification as Link Prediction on Multimodal Knowledge Graphs

### 31 Dec 2023 v1

@@ Abstract
Camera traps are valuable tools in animal ecology for biodiversity monitoring and conservation. However, challenges like poor generalization to deployment at new unseen locations limit their practical application. Images are naturally associated with heterogeneous forms of context possibly in different modalities. In this work, we leverage the structured context associated with the camera trap images to improve out-of-distribution generalization for the task of species identification in camera traps. For example, a photo of a wild animal may be associated with information about where and when it was taken, as well as structured biology knowledge about the animal species. While typically overlooked by existing work, bringing back such context offers several potential benefits for better image understanding, such as addressing data scarcity and enhancing generalization. However, effectively integrating such heterogeneous context into the visual domain is a challenging problem. To address this, we propose a novel framework that reformulates species classification as link prediction in a multimodal knowledge graph (KG). This framework seamlessly integrates various forms of multimodal context for visual recognition. We apply this framework for out-of-distribution species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets and achieve competitive performance with state-of-the-art approaches. Furthermore, our framework successfully incorporates biological taxonomy for improved generalization and enhances sample efficiency for recognizing under-represented species. [^1] Code and data will be released on GitHub.
@@ 1 Introduction
Towards this end, we propose a new learning framework, COSMO (Classification Of Species using Multimodal cOntext), where we first organize all species images and contextual information as a multimodal knowledge graph (KG) and then reformulate species classification as the standard link prediction task on the KG. Specifically, we consider species images, their corresponding labels (which are available in the training data), and their associated attributes provided in the context as entities within our KG (see Figure 1 for an example). We represent the relationships between these entities as edges in our KG (see a more concrete description of our KG construction in Section 3.2). Our KG is multimodal since its entities are from different modalities. In this context, species classification can be framed as a link prediction task, where the objective is to predict the presence of an edge between an image and its corresponding species label within the KG. This learning framework enables a unified way to incorporating heterogeneous contextual information for species classification. Each form of multimodal information is treated as a type of entity, a first-class citizen of the multimodal KG with its representation computed using a modality-specific encoder. The learning process enables the interaction of different modalities in a joint feature space for robust representation learning. In addition, COSMO demonstrates greater flexibility by not assuming uniform availability of all contextual information, unlike previous methods.
@@ 5 Results
    Does the use of multiple modalities as context contribute toward better performance? (Section 5.1)
@@ 5.2 Comparison with OOD Generalization Approaches
Model Prediction Taxonomy
COSMO (w/ taxonomy) ocelot (leopardus pardalis)
ocelot (leopardus pardalis) →  leopardus →  felinae →
felidae →  feliformia →  carnivora
COSMO (no-context) Gray fox (urocyon cinereoargenteus)
Gray fox (urocyon cinereoargenteus) →  urocyon →
canidae →  caniformia →  carnivora
COSMO (w/ taxonomy) Ocellated turkey (meleagris ocellata)
Ocellated turkey (meleagris ocellata) →  meleagris →
meleagridinae →  normal-⋯\;$\cdots$ →  birds (aves)
COSMO (no-context) Central American agouti (dasyprocta punctata)
Central American agouti (dasyprocta punctata) →
dasyprocta →  dasyproctidae →  hystricomorpha →  normal-⋯\;$\cdots$
→  mammals (mammalia)
Table 5: Comparison of COSMO model with and without taxonomy edges (iWildCam2020-WILDS validation set). The use of taxonomy information helps the model to avoid semantically implausible predictions.
@@ 5.4.1 Error Analysis for the Taxonomy-aware Model
We compare the predictions of COSMO which uses taxonomy to the no-context baseline (Table 5). Notably, the inclusion of taxonomy information assists the model in avoiding implausible predictions. For instance, consider the case of the animal ocelot (eopardus pardalis), which belongs to the cat family (feliformia). The use of taxonomy information prevents the misprediction of this animal as a gray fox, which belongs to the dog family (caniformia). Similarly, in the second example, the baseline model incorrectly predicts the given image as Central American agouti, a mammal, instead of ocellated turkey, a bird.
@@ 5.4.3 Performance Comparison for Under-represented Species
6 Conclusion
In this work, we presented a novel framework in which the species classification task is reformulated as link prediction in a multimodal KG of species images and their associated heterogeneous contexts. This enables a unified way to leverage various forms of multimodal context, e.g., numerical, categorical, and taxonomy information associated with images for species identification in camera traps. Through our experiments, we demonstrate that our framework achieves superior out-of-distribution (OOD) generalization and competitive performance with state-of-the-art for species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets. Additionally, our framework demonstrates improved sample efficiency in recognizing under-represented and vulnerable wildlife species. Future work will focus on integrating a more comprehensive spectrum of diverse contexts for use with real-world camera trap deployments.

### 08 Jan 2024 v2

@@ Abstract
Camera traps are valuable tools in animal ecology for biodiversity monitoring and conservation. However, challenges like poor generalization to deployment at new unseen locations limit their practical application. Images are naturally associated with heterogeneous forms of context possibly in different modalities. In this work, we leverage the structured context associated with the camera trap images to improve out-of-distribution generalization for the task of species identification in camera traps. For example, a photo of a wild animal may be associated with information about where and when it was taken, as well as structured biology knowledge about the animal species. While typically overlooked by existing work, bringing back such context offers several potential benefits for better image understanding, such as addressing data scarcity and enhancing generalization. However, effectively integrating such heterogeneous context into the visual domain is a challenging problem. To address this, we propose a novel framework that reformulates species classification as link prediction in a multimodal knowledge graph (KG). This framework seamlessly integrates various forms of multimodal context for visual recognition. We apply this framework for out-of-distribution species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets and achieve competitive performance with state-of-the-art approaches. Furthermore, our framework successfully incorporates biological taxonomy for improved generalization and enhances sample efficiency for recognizing under-represented species. [^1] Website: https://osu-nlp-group.github.io/COSMO/
@@ 1 Introduction
Towards this end, we propose a new learning framework, COSMO (Classification Of Species using Multimodal cOntext), where we first organize all species images and contextual information as a multimodal knowledge graph (KG) and then reformulate species classification as the standard link prediction task on the KG. Specifically, we consider species images, their corresponding labels (which are available in the training data), and their associated attributes provided in the context as entities within our KG (see Figure 1 for an example). We represent the relationships between these entities as edges in our KG (see a more concrete description of our KG construction in Section 3.2). Our KG is multimodal since its entities are from different modalities. In this context, species classification can be framed as a link prediction task, where the objective is to predict the presence of an edge between an image and its corresponding species label within the KG. This learning framework enables a unified way to incorporate heterogeneous contextual information for species classification. Each form of multimodal information is treated as a type of entity, a first-class citizen of the multimodal KG with its representation computed using a modality-specific encoder. The learning process enables the interaction of different modalities in a joint feature space for robust representation learning. In addition, COSMO demonstrates greater flexibility by not assuming uniform availability of all contextual information, unlike previous methods.
@@ 5 Results
    Does the use of contextual information contribute toward better performance? (Section 5.1)
@@
Figure 2: Comparison of COSMO model with and without taxonomy edges (iWildCam2020-WILDS validation set). The use of taxonomy information helps the model to avoid semantically implausible predictions.
@@ 5.4.1 Error Analysis for the Taxonomy-aware Model
We compare the predictions of COSMO which uses taxonomy to the no-context baseline (Figure 2). Notably, the inclusion of taxonomy information assists the model in avoiding implausible predictions. For instance, consider the case of the animal ocelot (eopardus pardalis), which belongs to the cat family (feliformia). The use of taxonomy information prevents the misprediction of this animal as a gray fox, which belongs to the dog family (caniformia). Similarly, in the second example, the baseline model incorrectly predicts the given image as Central American agouti, a mammal, instead of ocellated turkey, a bird.
@@ 5.4.3 Performance Comparison for Under-represented Species
6 Discussion and Conclusion
In this work, we presented a novel framework in which the species classification task is reformulated as link prediction in a multimodal KG of species images and their associated heterogeneous contexts. This enables a unified way to leverage various forms of multimodal context, e.g., numerical, categorical, and taxonomy information associated with images for species identification in camera traps. Through our experiments, we demonstrate that our framework achieves superior out-of-distribution (OOD) generalization and competitive performance with state-of-the-art for species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets. Additionally, our framework demonstrates improved sample efficiency in recognizing under-represented and vulnerable wildlife species.
We assume that there is a perfect linkage between these contexts and the corresponding images in the training set. However, in scenarios where such linkage is unavailable, the training procedure may introduce noise, which could lead to inferior generalization capabilities in the model. Additionally, it is important to note that the effectiveness of diverse contexts varies based on their informativeness for the given task. Interestingly, combining two or more contexts could degrade performance compared to using a single context type in some cases (Table 1). We hypothesize that specific metadata, like location, could exert a stronger regularization influence on improving generalization in species recognition tasks than other metadata. To improve this aspect, future work will involve enabling the model to assign higher importance to more informative metadata.
Furthermore, we are interested in training a foundation model for camera trap species identification across a wider spectrum of species. This model is expected to generalize more effectively to novel camera trap deployments across the world. Additionally, we would like to focus on integrating a more comprehensive spectrum of diverse contexts such as temperature, weather conditions, habitat, and sequence information for use with real-world camera trap deployments.